OpenStudy (cutiecomittee123):

how do find the vertex and foci of this hyperbola (x-2)^2/36 - (y+1)^2/64 =1

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

standard form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] you get the center with your eyeballs right?

OpenStudy (misty1212):

this is where you say "yes i see it" or "no i have no clue"

OpenStudy (cutiecomittee123):

yes:)

OpenStudy (misty1212):

ok what is the center?

OpenStudy (cutiecomittee123):

(2,-1)

OpenStudy (misty1212):

ok good now we need to know one other thing what does it look like left or right?